### v1

# SAGE: Sequence-level Adaptive Gradient Evolution for Generative Recommendation

Abstract.
While works such as OneRec have validated the scaling laws of Large Language Models (LLMs) in recommender systems, they rely on a cumbersome separate vocabulary. This dependency prevents the model architecture from reusing native LLM vocabularies, resulting in high maintenance costs and poor scalability. In response, we aim to efficiently reuse open-source LLM architectures without constructing a separate tokenization vocabulary. Furthermore, we identify that the optimization strategy of OneRec—Gradient-Bounded Policy Optimization (GBPO)—suffers from a ”Symmetric Conservatism” problem: its static gradient boundaries structurally suppress the update momentum required for cold-start items and fail to prevent diversity collapse in high-noise environments.To address this issue, we propose SAGE (Sequence-level Adaptive Gradient Evolution), a unified optimization framework tailored for list-wise generative recommendation. SAGE introduces two key innovations: (1) Sequence-level Signal Decoupling: By combining a geometric mean importance ratio with decoupled multi-objective advantages, we eliminate token-level variance and resolve the ”Reward Collapse” problem. (2) Asymmetric Adaptive Dynamics: We construct a dynamic gradient manifold that applies a ”Boost Factor” to high-potential cold-start items to achieve super-linear updates and employs an ”Entropy-Aware Penalty” to break information cocoons. Theoretical analysis and empirical results demonstrate that SAGE effectively unblocks cold-start traffic and sustains recommendation diversity, all while retaining the numerical stability of GBPO.
Generative Recommender Systems, Reinforcement Learning
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
In the digital age, recommender systems have become indispensable tools for user navigation of vast information landscapes. However, traditional methodologies, primarily leveraging historical interaction data and techniques such as collaborative filtering and matrix factorization, exhibit inherent shortcomings. Although effective in data-rich environments for capturing certain facets of user preference, these approaches often overlook valuable external knowledge, thereby contributing to the well-established ”filter bubble” effect (Koren et al., 2009). Moreover, industrial deployments frequently maintain a separation between model optimization and strategic iteration. This division, often manifested through distinct teams dedicated to model tuning versus strategy adjustment (Covington et al., 2016; Liu et al., 2017; Qin et al., 2022; Luo et al., 2024), can impede the rapid adaptation necessary to address evolving business requirements.
Recent advancements in large-scale pre-trained language models (LLMs) have introduced a transformative opportunity for the recommender system domain. These models demonstrate remarkable capabilities in semantic understanding and knowledge reasoning, and are capable of learning new tasks with limited data  (Brown et al., 2020; Devlin et al., 2019; Luo et al., 2024; Li et al., 2023; Yang et al., 2023; Liao et al., 2024; Ji et al., 2024; Li et al., 2025; Deng et al., 2025). Capitalizing on this potential to reshape data construction and model architecture in recommender systems holds the promise of overcoming the inherent limitations of traditional approaches, leading to enhancements in both the accuracy and diversity of personalized recommendations.
In generative recommendation systems such as OneRec, adapting an LLM to the recommendation task typically requires constructing a dedicated vocabulary to convert videos into Semantic IDs. Although this design alleviates the issues of unbounded vocabulary growth and cold-start problems inherent in traditional ID-based recommendation, it still faces several practical challenges.
First, semantic collision is a major issue: multiple distinct videos may be mapped to the exact same token sequence, while a large portion of tokens remain unmapped. This can lead to collision rates exceeding 30%, meaning that the model is unable to distinguish 30% of the items, resulting in severe information loss  (Zhou et al., 2025a). Second, in industrial-scale scenarios where hundreds of millions of new videos are uploaded daily, maintaining the tokenizer incurs substantial operational overhead.
Moreover, LLM adaptation training is mandatory. A dedicated training stage is required for the LLM to understand item tokens and absorb collaborative signals. General-purpose LLMs cannot interpret these newly defined “Itemic Tokens” (video IDs); thus, vocabulary expansion and alignment must be performed during the pre-training stage  (Zhou et al., 2025b, a). This includes initializing embeddings for the new tokens and freezing other parameters during training so that the LLM can learn the semantics of video tokens.
To circumvent these issues, we explore text-based models that do not require constructing a separate tokenization vocabulary, enabling efficient reuse of open-source LLM architectures. The approach is outlined as follows:
Data Construction Optimized for Large Language Models.
The core advantage of this paradigm lies in the fact that it can directly adopt the architecture of large language models without the need to construct a separate tokenization vocabulary, thus enabling the most efficient reuse of open-source large language models. User-side: We integrate user profiles with historical behavior information, constructing prompt inputs that combine static user attributes and dynamic interaction data. This ensures that the large model can capture both the basic characteristics and behavioral habits of the user. Item-side: We comprehensively leverage item descriptions, titles, and content-derived tags to provide a multi-faceted semantic representation of the items. Furthermore, we generate positive samples by constructing an item sequence based on the user’s subsequent interactions, guided by these tags. Label Construction: We generate the training label, an item sequence that accurately reflects the user’s preferences, based on their subsequent behavior data. This provides a more precise supervisory signal for model training.
Two-Stage Interactive Training Strategy.
SFT Stage: In the Supervised Fine-Tuning (SFT) stage, we employ a distilled version of DeepSeek R1 to fine-tune the large language model using the constructed prompt data derived from recommender system user interactions. This process activates the model’s inherent capabilities for recommendation tasks, facilitating knowledge transfer and initial model adaptation.Additionally, high-quality samples generated in the RLHF stage will be added to the training samples. RLHF Stage: Building upon the SFT stage, we propose SAGE (Sequence-level Adaptive Gradient Evolution), a unified optimization framework tailored for list-wise generative recommendation. SAGE introduces two key innovations: (1) Sequence-level Signal Decoupling: We propose a geometric mean importance ratio combined with a decoupled multi-objective advantage estimator. This eliminates token-level variance and resolves ”Reward Collapse” in multi-objective scenarios (e.g., balancing clicks and duration). (2) Asymmetric Adaptive Dynamics: SAGE constructs a dynamic gradient manifold that is asymmetric by design. It introduces a ”Boost Factor” to allow super-linear updates for high-potential cold-start items, and an ”Entropy-Aware Penalty” that dynamically intensifies gradients for low-diversity negative samples to break information cocoons. Theoretical analysis and empirical results demonstrate that SAGE effectively unblocks cold-start traffic and sustains recommendation diversity, all while retaining the numerical stability of GBPO.
Figure 1. Illustrates the motivation behind our research
Overall, Figure 1 illustrates the motivation behind our research. Our approach offers a novel perspective on recommender system design by seamlessly integrating the strengths of large language models and reinforcement learning. It effectively addresses the limitations of traditional methods—particularly the issue of information silos—while providing a solid technical foundation for multi-objective and customized recommendation in real-world business scenarios.
2. Related Work
2.1. Generative Recommendation and Its Challenges
With the breakthroughs in generative artificial intelligence, the paradigm of recommender systems is shifting from traditional Discriminative Ranking to Generative Recommendation. The core idea of generative recommendation is to reformulate the task as an end-to-end autoregressive generation problem, directly generating the Semantic ID sequence of the target item, thereby enabling direct optimization toward the final objective. This requires a tokenization strategy that maps items into discrete token sequences (Semantic IDs). OneRec is a pioneering paradigm in this direction, relying on predefined structured vocabularies  (Li et al., 2025; Deng et al., 2025).
However, this paradigm introduces a dependency on a heavy vocabulary: the model must memorize the mapping between semantic tokens and specific items. The discrete nature of these tokens makes optimization difficult, as even slight deviations in token generation may lead to valid but irrelevant item IDs, or even entirely invalid sequences  (Deng et al., 2025; Zhou et al., 2025b).
2.2. Evolution of Policy Optimization
To align large models with complex business objectives, reinforcement learning–based methods have become mainstream. GRPO  (Shao et al., 2024) eliminates the need for a Critic network by normalizing rewards within sampled groups, and it has been widely applied in reasoning tasks. However, GRPO suffers from the problem of Reward Collapse in multi-objective optimization, where samples with different performance may end up with identical advantage values after normalization, reducing the resolution of training signals. In addition, negative feedback dominates recommendation scenarios, and directly applying standard RL often leads to gradient instability  (Zhou et al., 2025b).
To address these issues, OneRec-V2  (Zhou et al., 2025b) proposed GBPO (Gradient-Bounded Policy Optimization). GBPO introduces a dynamic gradient boundary derived from binary cross-entropy (BCE), effectively preventing gradient explosion caused by negative samples and improving stability in large-scale training.
Although GBPO resolves the stability issue, its symmetric and conservative design ($r\leq 1$) introduces new bottlenecks. Cold-start video views dropped by 44.7%, because the strict upper bound on gradients prevents strong updates on sparse positive signals for new items. Cluster density increased by 11.7%, as the static boundary fails to preserve entropy-increasing exploratory behavior.
Therefore, we propose SAGE, a unified optimizer that is sequence-level, adaptive, and multi-objective.Token-level optimization introduces noise in Semantic ID generation. Following the logic of GSPO  (Zheng et al., 2025), SAGE elevates the importance ratio to the sequence level (geometric mean), ensuring the overall consistency of tokenized item IDs. BAPO  (Xi et al., 2025) shows that standard clipping suppresses exploration. SAGE adopts BAPO’s Entropy-Clip Rule but transforms it into a dynamic penalty term for negative samples, countering the diversity collapse observed in OneRec-V2. To offset GBPO’s suppression of new items, SAGE incorporates the Clip-Higher philosophy from DAPO  (Yu et al., 2025), enabling super-linear updates for positive cold-start signals. To address the reward collapse issue in GRPO, SAGE integrates the decoupled normalization strategy from GDPO  (Liu et al., 2026), achieving precise alignment across multiple objectives (e.g., clicks and watch time).
3. Method
The proposed recommender system framework comprises three primary components: Data Construction, the SFT Stage, and the RLHF Stage, as depicted in Figure  2.
Figure 2. An overview of the modeling approach used in RecLLM-R1.
3.1. Data Construction
To align with the input requirements of large language models, we constructed a training dataset specifically tailored for recommender systems by transforming features into natural language descriptions. User-side Data:We transformed user profiles (e.g., age, gender, interests) and recent interaction history (e.g., sequences of clicked, viewed, liked, bookmarked, or purchased items) into chronologically ordered textual descriptions, aggregated by topic. For instance, ”Over the past 7 days, the user viewed ’Sichuan Cuisine’ videos twice and bookmarked one article titled ’Chengdu Travel Guide’.” Item-side Data: We primarily utilized item descriptions, titles, and granular tag labels derived through content understanding, alongside statistical item features, to generate natural language descriptions. Label Construction: Based on user engagement metrics such as browsing duration and interaction behaviors (e.g., bookmarking, sharing, commenting, following, liking), we constructed a sequence of items that reflects the user’s eventual preferences as positive training samples. This ensures the authenticity and business relevance of the labels. Prompt Construction: We concatenated the user-side and item-side textual information. The objective is to explicitly recommend the most relevant sequence of items to the user, with candidate items drawn from the provided data pool.
Through this methodology, we transformed traditional structured and semi-structured recommendation data into a natural language format interpretable and processable by large language models (LLMs). This data construction approach not only preserves the benefits of traditional interaction data but also imbues large language models with rich semantic context, thereby mitigating the limitations of the ”filter bubble” effect.
3.2. SFT Stage: Activating the Recommender Capability of the Large Language Model
Following the construction of the textualized recommendation dataset, we proceeded to the Supervised Fine-Tuning (SFT) stage. The primary objective of this phase was not to immediately achieve optimal recommendation performance, but rather to ”activate” or ”cold-start” the pre-trained large language model (LLM), endowing it with the foundational ability to address recommendation tasks.
For efficiency, we used the pre-trained Qwen3-8B model and adopted the previously constructed dataset that is compact in scale yet high in quality, which includes high-quality sample data generated in the RLHF stage. In this stage, the emphasis was on leveraging a minimal amount of data to rapidly unlock the LLM’s potential within the recommendation domain, enabling it to learn the requisite input-output formats and develop a rudimentary understanding of recommendation task modeling. This process prepares the model for the subsequent, more intricate multi-step optimization.
Essentially, the SFT stage serves as a ”warm-up” providing the model with baseline recommendation capabilities prior to reinforcement learning. While SFT can already address certain recommendation needs—particularly cold-start scenarios, owing to the LLM’s general knowledge—it is inherently insufficient for optimizing complex long-term objectives, balancing multiple recommendation facets (such as diversity and novelty), or executing advanced reasoning. Therefore, SFT acts as a critical precursor, priming the model for comprehensive optimization in the subsequent RLHF stage. This design closely mirrors the training paradigm of DeepSeek R1.
3.3. RLHF Stage: Preference Alignment with Real-World User Interactions
After establishing foundational capabilities in the Supervised Fine-Tuning (SFT) stage, we introduce the RLHF phase of model. To address the cold-start and diversity issues mentioned in OneRec-V2, we propose SAGE (Sequence-level Adaptive Gradient Evolution). The core contribution of SAGE is the Asymmetric Dynamic Bounding mechanism, which adjusts the optimization trajectory based on the nature of the feedback (Positive vs. Negative) and the entropy of the generated list.
3.3.1. Adaptive Gradient Evolution
SAGE Objective Function. To address the scale inconsistency between the geometric mean ratio (token-scale) and the sum of log-probabilities (sequence-scale), we introduce a length-normalization term ($1/L$):
(1) $\begin{split}\nabla\mathcal{J}_{SAGE}(\theta)&=\mathbb{E}_{\tau\sim\pi_{old}}\Bigg[\frac{r_{slate}(\theta)}{\Phi(S,A)}\cdot A_{SAGE}\\ &\cdot\underbrace{\frac{1}{L}\sum_{t=1}^{L}\nabla\log\pi_{\theta}(i_{t}|u,S_{<t})}_{\text{Item-Averaged Gradient}}\Bigg]\end{split}$
  • $\frac{1}{L}$: Length Normalization Factor. This ensures the gradient magnitude aligns with the scale of $r_{slate}$ ($\approx 1$), preventing the optimizer from being biased towards longer sequences (though $L$ is fixed here, this maintains mathematical rigor).
  • $\Phi(S,A)$: The Adaptive Dynamic Denominator, defined below.
Case A: Positive Boost for Cold-Start (Solving Suppression). When the slate receives positive feedback ($A_{SAGE}\geq 0$), standard GBPO restricts the update ratio to $r\leq 1$. This is detrimental for cold-start items where the initial probability $\pi_{old}$ is extremely low. SAGE introduces a Boost Factor:
(2) $\Phi_{pos}=\begin{cases}\frac{1}{1+\epsilon_{boost}}&\text{if }r_{slate}(\theta)>1+\epsilon_{boost}\\ 1&\text{otherwise}\end{cases}$
  • $\epsilon_{boost}$: The Positive Boost Factor (e.g., $0.2\sim 0.5$).
  • Mechanism: If the model successfully identifies a high-potential item (resulting in a high ratio $r_{slate}$), SAGE locks the denominator such that the effective gradient coefficient is maintained at $1+\epsilon_{boost}$.
  • Impact: This allows the probability of successful cold-start items to grow super-linearly, providing the necessary momentum to break through the initial exposure threshold.
Case B: Diversity-Aware Penalty (Solving Collapse). When the slate receives negative feedback ($A_{SAGE}<0$), we distinguish between ”failed exploration” and ”homogenized repetition.” We utilize the List Entropy $H(S)$ to measure the semantic diversity of the slate.
(3) $\Phi_{neg}=\max\left(1,\frac{1-r_{slate}(\theta)}{\lambda(H(S))}\right)$
The regulation factor $\lambda$ is negatively correlated with diversity:
(4) $\begin{split}\lambda(H(S))=1+\beta\cdot\\ \tanh\left(\max(0,\bar{H}-H(S))\right)\end{split}$
  • $H(S)$: The content diversity score of the current slate (e.g., category entropy).
  • $\bar{H}$: The moving average of historical diversity.
  • $\beta$: The Diversity Temperature.
  • Mechanism:
    • Cocoon Penalty ($H(S)<\bar{H}$) If the slate is homogenized (low entropy) and rejected, $\lambda>1$. This shrinks the denominator $\Phi_{neg}$, resulting in a larger gradient (stronger penalty). This forces the model to escape the local optimum of recommending safe but boring items.
    • Exploration Tolerance ($H(S)\geq\bar{H}$) If the slate is diverse (high entropy) but rejected, $\lambda\approx 1$. The boundary reverts to the standard GBPO form, treating the error as a standard training sample without excessive punishment.
We illustrate the Comparison of Gradient Boundaries in  fig. 3 and the Evolution of Training Dynamics in  fig. 4.
Figure 3. Comparison of Gradient Boundaries. This figure illustrates the effective gradient coefficient ($r_{slate}/\Phi$) as a function of the importance ratio ($r_{slate}$). GBPO (Dashed Red) applies a static symmetric bound, capping positive updates at 1.0. In contrast, SAGE (Solid Blue) employs an asymmetric dynamic bound: it boosts high-potential positive samples beyond 1.0 (Case A) and applies entropy-aware penalties to negative samples (Case B), imposing stronger penalties on low-diversity slates (light blue) compared to high-diversity ones (dark blue).
Figure 4. Evolution of Training Dynamics. (a) SAGE rapidly boosts the generation probability of cold-start items compared to GBPO, demonstrating the effectiveness of the Positive Boost mechanism. (b) SAGE maintains consistently higher list diversity throughout training, verifying that the Entropy-Aware Penalty successfully prevents mode collapse, whereas GBPO tends to converge towards lower-entropy states.
3.3.2. Multi-Objective Signal Decoupling
To construct a high-fidelity learning signal from complex user feedback (clicks, duration, comments), we adopt a Decouple-then-Aggregate strategy inspired by GDPO.
Decoupled Slate Advantage Directly summing weighted rewards leads to ”Reward Collapse,” where distinct behaviors map to identical advantage values. Instead, we normalize each objective independently within the sampling group before aggregation:
(5) $\begin{split}A_{SAGE}(S)=\text{Norm}_{batch}\Bigg(\sum_{m=1}^{M}w_{m}\\ \cdot\frac{R_{m}(S)-\mu_{m}^{group}}{\sigma_{m}^{group}+\epsilon}\Bigg)\end{split}$
  • $R_{m}(S)$: The raw reward value for the $m$-th objective (e.g., watch time).
  • $\mu_{m}^{group},\sigma_{m}^{group}$: The mean and standard deviation of rewards within the sampled group $G$ for a single user query.
  • $\text{Norm}_{batch}$: A final batch-level normalization ($z$-score) to ensure $A_{SAGE}$ follows a standard normal distribution, providing a stable scale for the gradient boundaries.
Sequence-level Ratio Formulation Since user feedback is directed at the item level, token-level importance ratios introduce high variance. We define the importance ratio as the Geometric Mean of the token probability ratios over the generated list:
(6) $\begin{split}r_{slate}(\theta)=\exp\Bigg(\frac{1}{L}\sum_{t=1}^{L}\log\\ \frac{\pi_{\theta}(i_{t}|u,S_{<t})}{\pi_{\theta_{old}}(i_{t}|u,S_{<t})}\Bigg)\end{split}$
  • $L$: The fixed length of the recommendation slate ($L=6$).
  • Physical Meaning: This metric smooths out intra-item token fluctuations, representing the average probability gain of generating the entire slate under the new policy versus the old policy.
4. Experiments
In this section, we conduct comprehensive experiments to evaluate the effectiveness of the proposed framework. We aim to validate whether our Native-LLM + SAGE paradigm (”RecLLM”) can outperform state-of-the-art generative baselines in terms of accuracy, while simultaneously resolving the cold-start and diversity bottlenecks inherent in the OneRec-V2 paradigm.
4.1. Experimental Setup
4.1.1. Datasets
Public Datasets: We evaluated the proposed framework on three publicly available real-world benchmarks from the Amazon Product Reviews dataset  (He and McAuley, 2016), which encompasses user reviews and item metadata spanning from May 1996 to July 2014. Specifically, we utilized three categories of this dataset for the sequential recommendation task: “Sports” “Beauty” and “Toys”.
4.1.2. Baselines
We compare our method against distinct categories of baselines:
  • SASRec: The classic self-attentive sequential model.
  • Generative/Retrieval Models: ReaRec, TIGER, and LC-Rec, which utilize VQ-VAE or retrieval-augmented generation.
  • OneRec Variants:
    • OneRec-V2: The SOTA generative baseline using Semantic IDs and GBPO optimizer.
    • OneRec-SAGE: Applying our SAGE optimizer to the traditional OneRec (Semantic ID) architecture to test optimizer effectiveness in isolation.
    • Only SFT (RecLLM): The Native-LLM model trained only with Supervised Fine-Tuning.
    • RecLLM: The proposed framework combining the Native-LLM architecture with the SAGE optimizer.
4.1.3. Evaluation Metrics
We evaluate performance across three dimensions using top-$K$ metrics (default $K=10$):
  • Accuracy: Normalized Discounted Cumulative Gain (NDCG@K) and Recall (Recall@K).
  • Diversity: Intra-List Diversity (ILD), measuring the average semantic distance between items within a generated slate, and Entropy@K, measuring the distribution of recommended categories across the test set.Since all items in a dataset category belong to the same root (e.g., ”Beauty”), calculating entropy at the root level is trivial. Therefore, we utilize the second-level sub-categories (e.g., ”Skin Care”, ”Makeup”, ”Fragrance” for Beauty) provided in the item metadata to calculate Shannon Entropy. This measures whether the model recommends a diverse range of granular interests rather than converging to a single sub-niche.
  • Cold-Start: Given that precise item publication timestamps are unavailable in this dataset, we adopt the standard academic definition for cold-start evaluation. We define Cold Items as the set of items belonging to the bottom 20% of the interaction frequency distribution in the training set. This metric evaluates the model’s ability to discover and recommend long-tail items that have sparse historical interactions, preventing the ”rich-get-richer” feedback loop.
4.1.4. Implementation Details
The base model is Qwen3-8B. We set the slate length $L=6$. The SFT stage is trained for 1 epochs. In the RLHF stage, we use a learning rate of 5e65e-6 with the AdamW optimizer. For SAGE specific hyperparameters, based on validation set tuning, we set the Positive Boost Factor $\epsilon_{boost}=0.3$, the Diversity Temperature $\beta=0.5$.
4.2. Main Results
Superiority of RecLLM: As shown in Table  1, our proposed method (”RecLLM”) consistently outperforms all baselines across all datasets. Specifically, compared to the strongest baseline OneRec-V2, RecLLM achieves improvements of 4.3% to 8.9% in NDCG metrics. This confirms that directly reusing the Native LLM vocabulary—when optimized correctly—can surpass models relying on heavy, separate tokenization.
The Necessity of SAGE: Notably, the ”Only SFT” model performs poorly, significantly lagging behind even SASRec. SFT mainly functions to activate the model’s fundamental capabilities. without reinforcement learning to align the Native LLM with recommendation goals, the model fails to capture complex user preferences. Furthermore, OneRec-SAGE (applying SAGE to the old architecture) shows mixed results: while competitive, it slightly underperforms OneRec-V2 in some metrics (e.g., R@5 on Beauty). This suggests that SAGE’s aggressive exploration requires the robust semantic representation of the RecLLM to fully realize its potential without destabilizing the training.
Table  2 investigates the impact of optimization strategies on recommendation diversity (Entropy) and long-tail item discovery (Cold-Recall).
Breaking the Symmetric Conservatism: Comparing RecLLM against OneRec-V2, we observe substantial gains:
  • Cold-Recall: SAGE improves cold-start performance by 49% (Beauty), 63% (Sports) and 43% (Toys). This validates that the Positive Boost Factor ($\epsilon_{boost}$) effectively allows high-potential fresh items to break through the symmetric gradient barrier of GBPO.
  • Diversity: Entropy improves by 5.9% , 10.36% and 7.36%, confirming that the Entropy-Aware Penalty prevents the model from collapsing into safe, popular items.
The Accuracy-Diversity Trade-off: It is worth noting that OneRec-SAGE achieves the highest absolute values for Entropy and Cold-Recall. However, revisiting Table  1, we see that OneRec-SAGE suffers a drop in top-k accuracy (e.g., lower R@5 than OneRec-V2). This indicates that applying aggressive exploration on the limited semantic ID vocabulary can lead to over-exploration. In contrast, RecLLM achieves the most Optimal state: it delivers the highest overall accuracy table 1 while still providing a massive upgrade in diversity and cold-start handling compared to the current SOTA baseline (OneRec-V2).
Table 1. Overall performance comparison on three Amazon Product Reviews datasets. Bold indicates the best performance, and underline indicates the second best.
Dataset Beauty Sports Toys
Method R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
SASRec 0.0393 0.0639 0.0209 0.289 0.0240 0.0389 0.0130 0.0178 0.0420 0.0658 0.0217 0.0294
ReaRec 0.0488 0.0702 0.0341 0.0231 0.0348 0.0152 0.0189 0.0249 0.0517 0.0706 0.0369 0.0430
TIGER 0.0413 0.0628 0.0277 0.0346 0.0216 0.0331 0.0145 0.0182 0.0367 0.0527 0.0255 0.0307
LC-Rec 0.0495 0.0764 0.0338 0.0424 0.0269 0.0418 0.0177 0.0225 0.0350 0.0549 0.0221 0.0285
OneRec-V2 0.0646 0.0929 0.0456 0.0550 0.0365 0.0551 0.0252 0.0317 0.0693 0.0958 0.0496 0.0581
OneRec-SAGE 0.0612 0.0931 0.0424 0.0557 0.0351 0.0541 0.0246 0.0297 0.0671 0.0960 0.0466 0.0583
Only SFT (RecLLM) 0.0291 0.023 0.0158 0.204 0.0163 0.0263 0.0104 0.0101 0.0386 0.0571 0.0169 0.0212
RecLLM 0.0683 0.0973 0.0497 0.0590 0.0385 0.0575 0.0271 0.0339 0.0738 0.1018 0.0533 0.0622
Improve +5.67% +4.7% +8.9% +7.2% +5.46% +4.33% +7.35% +6.96% +6.5% +6.3% +7.48% +7.11%
Table 2. Relative to OneRec‑v2, SAGE delivers significant gains in cold‑start recovery and recommendation diversity.
Dataset Beauty Sports Toys
Method Entropy@10 Cold-Recall Entropy@10 Cold-Recall Entropy@10 Cold-Recall
OneRec-V2 2.285 0.0326 2.274 0.0282 2.315 0.0365
OneRec-SAGE 2.551 0.0619 2.528 0.0713 2.574 0.0695
RecLLM 2.420 0.0487 2.510 0.0460 2.485 0.0522
Improve +5.9% +49% +10.36% +63% +7.36% +43%
4.3. Ablation Study
To understand the contribution of each component, we conduct an ablation study on the Amazon-Beauty dataset. The results are visually presented in Figure  5.
Figure 5. Ablation Study on Amazon-Beauty Dataset. Metrics are normalized to the full SAGE model. Removing individual components leads to degradation in their respective targeted areas (e.g., removing Boost hurts Cold-Start most), validating the design of each module.
  • w/o Positive Boost: Setting $\epsilon_{boost}=0$ reverts the positive boundary to standard GBPO ($r\leq 1$). We observe a drastic drop in Cold-Start Recall (Green Bar) to 0.50 of the full model’s performance, confirming this component’s critical role in item exploration.
  • w/o Entropy Penalty: Removing the entropy term and using the static GBPO boundary for negative feedback leads to a significant decrease in Entropy@10 (Orange Bar) to 0.78, indicating a collapse towards redundant recommendations.
  • w/o Signal Decoupling: Using a simple weighted sum of rewards instead of decoupled normalization results in a noticeable drop in NDCG@10 (Blue Bar) to 0.90 and a decrease in other metrics. This suggests that ”Reward Collapse” was indeed hindering the precise multi-objective optimization.
The full SAGE model achieves the best balance, demonstrating that the asymmetric adaptive dynamics and signal decoupling are complementary and essential.
5. Conclusion
In this paper, we identified fundamental limitations in existing optimization strategies for generative recommendation, specifically the ”Symmetric Conservatism” of Gradient-Bounded Policy Optimization (GBPO), which leads to cold-start suppression and diversity collapse. To address these challenges within a list-wise generation paradigm, we proposed SAGE (Sequence-level Adaptive Gradient Evolution).
SAGE introduces a novel theoretical framework featuring **Asymmetric Adaptive Dynamics**. By applying a ”Positive Boost” to high-potential successful slates and an ”Entropy-Aware Penalty” to low-diversity failures, SAGE dynamically aligns the optimization trajectory with the long-tailed, high-noise nature of recommendation feedback. Furthermore, we incorporated Multi-Objective Signal Decoupling to ensure high-fidelity gradient signals in complex real-world scenarios.
RecLLM demonstrates that native‑vocabulary reuse not only surpasses models dependent on large standalone tokenizers but also maximizes the utility of the underlying LLM. While the absence of an independent vocabulary necessitates language‑level item encoding—posing challenges for real‑time systems—this constraint is expected to fade with increasing computational capacity.